Two cars leave an intersection. One car travels north; the other travels east.
When the car traveling north had gone 24 miles, the distance between the cars
was four miles more than three times the distance traveled by the car heading
east. Find the distance between the cars at that time.

Answer by mananth(16946)

Two cars leave an intersection. One car travels north; the other travels east.
When the car traveling north had gone 24 miles, the distance between the cars
was four miles more than three times the distance traveled by the car heading
east. Find the distance between the cars at that time.
..
let distance of car going east be x
by then car going north has traveled 24 miles
..
3x-4 is the distance between the two cars.
they form a right angle with the intersection point
..
(3x-4)^2=x^2+24^2
9x^2-24x+16=x^2+576
9x^2-x^2-24x-576+16=0
8x^2-24x-560=0
/4
2x^2-6x-140=0
2x^2-20x+14x-140=0
2x(x-10)+14(x-10)=0
(2x+14)(x-10)=0
x= 10 miles
..
m.ananth@hotmail.ca